What is the major and minor axis of an ellipse given the equation below:
16x^2 + 9y^2 -144 = 0
What is the location of the foci?
Thank you!

Answer by lwsshak3(11628)

What is the major and minor axis of an ellipse given the equation below:
16x^2 + 9y^2 -144 = 0
What is the location of the foci?
**
16x^2 + 9y^2 -144 = 0
divide by 144
x^2/9 + y^2/16 -1 = 0
x^2/9 + y^2/16 =1
This is an equation of an ellipse with vertical major axis.
Its standard form: (x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k)=(x,y) coordinates of center
For given equation:
center:(0,0)
a^2=16
a=√16=4
length of vertical major axis=2a=8
..
b^2=9
b=√9=3
length of minor axis=2b=6
..
c^2=a^2-b^2=16-9=7
c=√7
foci: (0,ąc+0)= (0,-√7) and(0,√7)